Stosch Company's balance sheet reported assets of $50,000, liabilities of $25,000 and common stock of $5,000 as of December 31, Year 1. If Retained Earnings on the balance sheet as of December 31, Year 2, amount to $18,000 and Stosch paid a $14,000 dividend during Year 2, then the amount of net income for Year 2 was which of the following?

A) $17,000

B) $19,000

C) $12,000

D) $21,000
Answer #1

Total assets=Total liabilities+Total equity

50,000=25,000+(Common stock+Retained earnings)

50,000=25,000+(5000+Retained earnings)

Retained earnings=50,000-25,000-5,000

=$20,000

Ending Retained earnings=Beginning Retained earnings+Net income-Dividends

18,000=20,000+Net income-14,000

Net income=18,000+14,000-20,000

=$12,000